Required information (The following information applies to the questions displayed below.] Sweeten Company had no jobs in pro3. What was the total manufacturing cost assigned to Job P
Answer #1

Solution:

Total estimated Overhead = (2500*1.40) + (1500*2.20) + $25000 = $31,800

Predetermined Overhead rate = $31800/ 4000 = $7.95 per Machine hour

Overhead assigned to Job P = 2300* $7.95 = $18,285

Total manufacturing cost assigned to Job P = 13000 + 21000+ 18285 = $52,285